OpenStudy (anonymous):

How do you determine the vertex of this equation f(x)= - 1/2x^2-x+3/2. By completing the square? THanks.

OpenStudy (anonymous):

yes. Completing the square will put the equation into vertex form \[y = m(x-p)^2 + q\] the vertex will then be (p,q).

OpenStudy (anonymous):

How do you complete the squre should of have been the question.

OpenStudy (anonymous):

\[ y=-\frac{1}{2}x^2-x + \frac{3}{2}\] becomes \[-2y=x^2 +2x -3\] becomes \[-2y = x^2 +2x +1^2 -1^2 -3\] \[-2y = \left[x^2 +2x +1^2\right] -1^2 -3\] becomes \[-2y = \left[(x+1)^2\right] -4\] becomes \[y = -\frac{1}{2}(x+1)^2 +2\] p=-1, q=2 vertex =(-1,2)

OpenStudy (anonymous):

Thank you very much!